OpenStudy (fantalux):

The perimeter of a rectangle is 96 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 7: 5.

OpenStudy (fantalux):

@sangya21 :(

OpenStudy (fantalux):

oh ur back...lol

OpenStudy (anonymous):

length/width =7/5 l = 7/5w P = 2(l+w) 96 = 2(l +7/5 l)

OpenStudy (fantalux):

ty!

OpenStudy (gschibby):

Well, here you get two equations:\[\frac{ length(x) }{ width(y) }=\frac{ 7 }{ 5 }\]and\[2x+2y=96\]

OpenStudy (fantalux):

30 ft by 18 ft 28 ft by 20 ft 24 ft by 24 ft 22 ft by 26 ft

OpenStudy (fantalux):

those are the possibilities...

OpenStudy (anonymous):

anytime :)

OpenStudy (fantalux):

@sangya21 so which one of those is it?

OpenStudy (anonymous):

length/width =7/5 l = 7/5w P = 2(l+w) 96 = 2(w+7/5 w) 96/2 = (5w +7w)/5 48*5 = 12w 240 = 12 w w = 20 l = 7/ 5* w l = 7/5*20 l = 7*4 = 28

OpenStudy (fantalux):

...??? plz just say which of these it is 30 ft by 18 ft 28 ft by 20 ft 24 ft by 24 ft 22 ft by 26 ft

OpenStudy (fantalux):

im really confused with your reply...

OpenStudy (anonymous):

length = 28 width = 20 so answer will be?

OpenStudy (fantalux):

thx. :)

OpenStudy (anonymous):

anytime. :)